# Refining Neural Activation Patterns for Layer-Level Concept Discovery in Neural Network-Based Receivers

### v1

@@ Abstract
Concept discovery in neural networks often targets individual neurons or human-interpretable features, overlooking distributed layer-wide patterns. We study the Neural Activation Pattern (NAP) methodology, which clusters full-layer activation distributions to identify such layer-level concepts. Applied to visual object recognition and radio receiver models, we propose improved normalization, distribution estimation, distance metrics, and varied cluster selection. In the radio receiver model, distinct concepts did not emerge; instead, a continuous activation manifold shaped by Signal-to-Noise Ratio (SNR) was observed—highlighting SNR as a key learned factor, consistent with classical receiver behavior and supporting physical plausibility. Our enhancements to NAP improved in-distribution vs. out-of-distribution separation, suggesting better generalization and indirectly validating clustering quality. These results underscore the importance of clustering design and activation manifolds in interpreting and troubleshooting neural network behavior.
@@ 1 Introduction
Neural networks are increasingly used in high-stakes applications (e.g., Jumper et al. (2021))—yet their internal representations remain difficult to interpret (Molnar (2022)). This opacity complicates debugging, trust, and accountability, especially in domains where model behavior must align with physical or safety constraints (Huyen (2022); Adebayo et al. (2020)). With regulatory frameworks like the EU AI Act mandating transparency in AI systems (European Commission (2024, 2019)), interpretable machine learning has become both a technical and societal necessity.
6G is a high-stakes application, projected to reach a \$68.7B global market by 2035 with a CAGR of 76.9% (MarketsAndMarkets (2024)), with the first commercial deployments expected by 2030 (Nokia (2025)). It promises extreme data rates, ultra-low latency, and dense connectivity, driving the need for advances in spectrum use, ultra-massive MIMO, and adaptive signal processing (Nokia Bell Labs (2024)). At the core is the physical layer, which sets the fundamental limits for capacity, efficiency, and reliability. Errors here propagate upward, making its performance critical. Traditional signal processing (Goldsmith (2005); Tse and Viswanath (2005)) is reaching its limits—motivating a shift to data-driven models with better adaptability (Hoydis et al. (2021)).
We explore this shift in the context of data-driven physical-layer receivers, where emerging 6G architectures are expected to replace hand-crafted pipelines with end-to-end learned models (Hoydis et al. (2021)). Architectures like DeepRx (Honkala et al. (2021)) process raw physical-layer radio signals to decode transmitted information under dynamic, noisy channel conditions. Consequently, these models introduce a new challenge: interpretability (Tuononen et al. (2025)). Physical-layer inputs are non-semantic, combining transmitted data with unpredictable environmental effects (e.g., noise, fading, interference), making most standard interpretability methods ineffective (Chergui et al. (2025); Fiandrino et al. (2022)).
To address this, we build on the Neural Activation Pattern (NAP) methodology (Bäuerle et al. (2022)), which clusters full-layer activation distributions to identify distributed, layer-level concepts. We extend NAP with improved normalization, distribution estimation, distance metrics, and varied cluster selection—as outlined in Figure 1—resulting in fewer unclustered samples and more stable clustering, and better out-of-distribution detection accuracy. To interpret the discovered clusters, we compare them with known physical parameters such as Signal-to-Noise Ratio (SNR). While no discrete cluster structure emerged, lower-dimensional embeddings revealed a continuous activation manifold aligned with SNR—suggesting neural receivers implicitly learn representations structured by core physical properties, even without explicit supervision.
@@ 2 Methodology
We refine the original NAP methodology (Bäuerle et al. (2022); detailed in Appendix A) to improve the stability and generalizability of the clustering results. In addition, we propose a principled approach to interpreting and evaluating clustering in the context of neural network-based systems, with a focus on radio receivers. Our refinements are shown in Figure 2 and described below:
$\phantom{,}\hat{a}^{l}_{i,u}=\frac{a^{l}_{i,u}-\mu^{l}_{u}}{\sigma^{l}_{u}},$ (1)
Proposed Location Disentanglement: We employ a form of non-parametric density estimation technique, known as Kernel Density Estimator (KDE; Rosenblatt (1956); Parzen (1962)), to aggregate the normalized activations per convolutional channel. For KDE, the estimated Probability Density Function (PDF) for channel $u$ in layer $l$ is given by

### v2

# Out-of-Distribution Detection via Channelwise Feature Aggregation in Neural Network–Based Receivers

@@ Abstract
Neural network–based radio receivers are expected to play a key role in future wireless systems, making reliable Out-Of-Distribution (OOD) detection essential. We propose a post-hoc, layerwise OOD framework based on channelwise feature aggregation that avoids classwise statistics–critical for multi-label soft-bit outputs with astronomically many classes. Receiver activations exhibit no discrete clusters but a smooth Signal-to-Noise-Ratio (SNR)–aligned manifold, consistent with classical receiver behavior and motivating manifold-aware OOD detection. We evaluate multiple OOD feature types, distance metrics, and methods across layers. Gaussian Mahalanobis with mean activations is the strongest single detector, earlier layers outperform later, and SNR/classifier fusions offer small, inconsistent AUROC gains. High-delay OOD is detected reliably, while high-speed remains challenging.
@@ 1 Introduction
Deep neural network–based radio receivers are emerging as a key component of future 6G systems hoydis2021toward, offering improved performance and reduced design complexity compared to conventional signal-processing pipelines farhadi2025. However, neural receivers can be brittle under distribution shift: when propagation conditions at inference deviate from those seen during training, performance may degrade sharply without any internal indication of failure luostari2025otavalidation. As emphasized in recent work on AI safety bengio2024extremerisks, monitoring and understanding the internal behavior of neural networks is essential for reliable deployment. In wireless systems, which operate under highly variable and unpredictable channel conditions, this requirement is fundamental.
Out-Of-Distribution (OOD) detection is therefore becoming indispensable in wireless machine learning liu2022oodwireless. We focus on Type-2 OOD (covariate shift), where the input distribution changes, while the prediction task remains unchanged, as in deployment scenarios with larger delay spreads, higher user velocities, or unforeseen interference conditions. Neural receivers such as DeepRx honkala2021deeprx produce thousands of soft-bit outputs per sample, forming a multi-label prediction problem tarekegn2024mll–implying an astronomical output space (Fig. 1, top). This makes classwise OOD detection methods fundamentally ill-posed, as they rely on a limited number of discrete classes.
Figure 1: Visual overview of OOD detection for neural receivers. OOD detection operates on aggregated activation features, rather than classwise statistics, forming an SNR-aligned manifold.
Practical deployment further constrains OOD detection: the receiver architecture is typically frozen and tightly optimized for latency, power consumption, and hardware compatibility tuononen2025interpreting. Many powerful OOD approaches–such as Bayesian neural networks blundell2015, deep ensembles lakshminarayanan2017, evidential networks aguilar2025evidential, or auxiliary OOD heads hendrycks2019–require architectural modification, retraining, or additional inference passes, making them unsuitable for production-grade wireless receivers. This motivates post-hoc OOD detection methods that operate solely on internal activations of a frozen model.
In this paper, we propose a post-hoc OOD detection framework for neural network–based radio receivers (Fig. 1, bottom) that avoids classwise modeling and leaves the receiver architecture unchanged. We show that receiver activations do not form discrete clusters but instead lie on a continuous Signal-to-Noise Ratio (SNR)–aligned manifold, a physically plausible structure that motivates manifold-aware OOD detection. Building on this insight, we develop a layerwise, channelwise feature aggregation framework that supports multiple OOD models and metrics, incorporates SNR-aware conditioning, and enables both intra- and inter-layer scoring. The approach enables both offline debugging and online monitoring, including anomaly detection yang2024oodsurvey, drift awareness uzlaner2025, and network-assisted reporting ren2021ood, contributing toward reliable and trustworthy neural receivers for future 6G systems.
2 Related Work
According to the manifold hypothesis, high-dimensional data governed by continuous generative factors tends to concentrate near low-dimensional manifolds embedded in the input space nilsson2007regressionmanifolds; bengio2013representation. Such continuous structures have been observed in domains ranging from images tenenbaum2000 to biomedical data islam2023, whereas discrete cluster structures are more typical of single-label classification tasks bengio2013representation.
Understanding internal neural representations is the goal of mechanistic interpretability, which studies how networks encode and manipulate information rauker2023survey; sharkey2025openproblems. Prior work spans visualization zeiler2014visualizing, concept-based probes kim2018, representation-similarity kornblith2019cka, and internal circuits or neuron communities wang2023circuit; lan2025inner. In wireless receivers, DeepRx has been analyzed through representation-level encoding of channel parameters such as SNR tuononen2025interpreting, underscoring the practical value of interpretability for deployed systems jeffares2025position.
OOD detection aims to identify inputs that deviate from the training distribution due to covariate or semantic shift yang2024oodsurvey. Many post-hoc methods operate on hidden representations, including $k$-Nearest Neighbors cover1967; papernot2018deepknn, class-conditional Mahalanobis lee2018mahalanobis; muller2025mahalanobis, and channel-selection for semantic shifts yuan2024discriminability. Extensions to multi-label settings include energy-based aggregation wang2021energy, which assumes semantic shift and conditional label independence zhang2023theoretical, and latent-variable models such as $\beta$-Variational Auto Encoder ($\beta$-VAE) for discrete generative factors sundar2020betavae. OOD score fusion methods include conformal p-value aggregation magesh2023conformalpvalues and one-class classifiers such as One-Class Support Vector Machine (OC-SVM scholkopf2001ocsvm; bounsiar2014ocsvm).
In wireless communications, OOD detection has mainly focused on transmitter identification and modulation recognition liu2022oodwireless. Related anomaly-detection work includes Deep Convolutional Autoencoder (DCAE)–based detection of physical tampering at Orthogonal Frequency-Division Multiplexing (OFDM) receivers dehmollaian2021dcae, where natural channel variation is excluded. Despite this extensive literature, post-hoc OOD detection for multi-label, Type-2 OOD neural physical-layer receivers–where classwise modeling is infeasible and the architecture must remain frozen–has not been previously addressed.
@@ 3 Methodology
Let $X\subset\mathbb{C}^{N_{f}\times N_{t}}$ denote the space of OFDM resource-grid 3gpp38211 inputs consisting of $N_{f}$ subcarriers and $N_{t}$ OFDM symbols (e.g., $192\times 14$). For a given modulation order $M$, the receiver outputs $Y\subset[0,1]^{N_{f}\times N_{t}\times B}$, where $B=\log_{2}M$ is the number of bits per symbol; for example, $B=4$ for 16-Quadrature Amplitude Modulation (16-QAM). We consider a frozen neural receiver $F_{\theta}$, parameterized by $\theta$, as a mapping $F_{\theta}:X\to Y$. During receiver training, data came from the joint distribution $P_{\text{train}}(X,Y)$, while deployment samples follow $P_{\text{test}}(X,Y)$. In Type-2 OOD detection (covariate shift quinonero2008datasetshift),
$P_{\text{train}}(Y\mid X)=P_{\text{test}}(Y\mid X)\text{, but}\,P_{\text{train}}(X)\neq P_{\text{test}}(X),$ (1)
as occurs when propagation characteristics–SNR, Doppler spread, multipath profile, interference–differ between training and deployment. Our task is to construct a detector $D$ that maps each received signal to a real-valued OOD score.
Raw physical- layer signal $X=\{x_{i}\}$ Frozen neural receiver $F_{\theta}(x_{i})$ Layerwise activations $A_{l}(x_{i})$ OOD score and decision $S_{i}\geq\tau$ OOD detector (e.g. k-NN) $S_{i}=D(g_{l,i})$ Channelwise aggregated + SNR features $g_{l,i}$ ; $p_{i}$
Figure 2: Post‑hoc OOD detection operating on a frozen model.
Layerwise Activation Extraction. Let the neural receiver contain $L$ convolutional layers. For a chosen layer $\ell\in\{1,\dots,L\}$, the activation tensor is
$A_{\ell}(x)\in\mathbb{R}^{C_{\ell}\times H_{\ell}\times W_{\ell}},$ (2)
where $C_{\ell}$ is the number of convolutional channels and ($H_{\ell}$$W_{\ell}$) are the spatial dimensions (i.e., $(N_{f},N_{t})$ for DeepRx). For a dataset $X=\{x_{i}\}_{i=1}^{N}$, we denote activations
$A_{\ell}(X)=\bigl\{A_{\ell}(x_{i})\bigr\}_{i=1}^{N}.$ (3)
Channelwise Feature Extraction. Features used for OOD detection are derived from per-channel activation vectors:
$\mathbf{a}_{\ell,i,c}=\mathrm{vec}\bigl(A_{\ell}(x_{i})[c,:,:]\bigr)\in\mathbb{R}^{H_{\ell}W_{\ell}},\,c=1,\dots,C_{\ell}.$ (4)
Each feature family applies a normalization scheme appropriate to the statistic being computed, and all resulting per-channel features are concatenated across channels. Thus, for each layer $\ell$ and input sample $x_{i}$, the activation-derived feature vector used for downstream modeling is one of
$\mathbf{g}_{\ell,i}\in\bigl\{\mathbf{m}_{\ell,i},\;\mathbf{k}_{\ell,i},\;\mathbf{e}_{\ell,i}\bigr\}.$ (5)
(i) Mean-activation features.
The MaxAbs scaled mean activation for channel $c$ is computed as  bauerle2022nap
$m_{\ell,i,c}=\frac{1}{H_{\ell}W_{\ell}}\sum_{q=1}^{H_{\ell}W_{\ell}}\frac{\mathbf{a}_{\ell,i,c}}{\max_{j}|\mathbf{a}_{\ell,j,c}|}[q],$ (6)
where $q$ indexes the flattened spatial locations. Concatenating across channels gives the mean feature vector
$\mathbf{m}_{\ell,i}=[m_{\ell,i,1},\dots,m_{\ell,i,C_{\ell}}].$ (7)
(ii) Activation-distribution features.
The activation-distribution shape for channel $c$ is computed by first applying Z-score normalization han2012datamining,
$\tilde{\mathbf{a}}_{\ell,i,c}^{\mathrm{Zscore}}=\frac{\mathbf{a}_{\ell,i,c}-\mu_{\ell,c}}{\sigma_{\ell,c}},$ (8)
and then computing a one-dimensional Kernel Density Estimate (KDE) parzen1962,
$\hat{p}_{\ell,i,c}(t)=\frac{1}{H_{\ell}W_{\ell}\,h_{\ell,c}}\sum_{q=1}^{H_{\ell}W_{\ell}}K\!\left(\frac{t-\tilde{\mathbf{a}}_{\ell,i,c}^{\mathrm{Zscore}}[q]}{h_{\ell,c}}\right),$ (9)